$SPAD/src/input laplace.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool laplace.output
)set message test on
)set message auto off
)clear all

———-

Some laplace transforms — * —

--S 1 of 27
f n == t**(n-1)*exp(-a*t)/factorial(n-1)
--R
--R                                                                   Type: Void
--E 1

———-

L(f n) is 1/(s + a)^{n} — * —

--S 2 of 27
f 2
--R
--R   Compiling function f with type PositiveInteger -> Expression(Integer
--R      )
--R
--R            - a t
--R   (2)  t %e
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 27
laplace(%, t, s)
--R
--R
--R               1
--R   (3)  --------------
--R         2           2

--R        s  + 2a s + a
--R                                                    Type: Expression(Integer)
--E 3
--S 4 of 27
f 5
--R
--R
--R         4  - a t
--R        t %e
--R   (4)  ---------
--R            24
--R                                                    Type: Expression(Integer)
--E 4
--S 5 of 27
laplace(%, t, s)
--R
--R
--R                            1
--R   (5)  ----------------------------------------
--R         5       4      2 3      3 2     4     5
--R        s  + 5a s  + 10a s  + 10a s  + 5a s + a
--R                                                    Type: Expression(Integer)
--E 5
--S 6 of 27
sin(a*t) - a*t*cos(a*t)
--R
--R
--R   (6)  sin(a t) - a t cos(a t)
--R                                                    Type: Expression(Integer)
--E 6
--S 7 of 27
laplace(%, t, s)
--R
--R
--R                3
--R              2a
--R   (7)  ---------------
--R         4     2 2    4
--R        s  + 2a s  + a
--R                                                    Type: Expression(Integer)
--E 7
--S 8 of 27
(cosh(a*t) - cos(a*t))/(2*a**2)
--R
--R

--R        cosh(a t) - cos(a t)
--R   (8)  --------------------
--R                   2
--R                 2a
--R                                                    Type: Expression(Integer)
--E 8
--S 9 of 27
laplace(%, t, s)
--R
--R
--R           s
--R   (9)  -------
--R         4    4
--R        s  - a
--R                                                    Type: Expression(Integer)
--E 9
--S 10 of 27
exp(-a*t) * sin(b*t) / b**2
--R
--R
--R           - a t
--R         %e     sin(b t)
--R   (10)  ---------------
--R                 2
--R                b
--R                                                    Type: Expression(Integer)
--E 10
--S 11 of 27
laplace(%, t, s)
--R
--R
--R                     1
--R   (11)  ------------------------
--R            2             3    2
--R         b s  + 2a b s + b  + a b
--R                                                    Type: Expression(Integer)
--E 11
--S 12 of 27
sin(a*t) * cosh(a*t) - cos(a*t) * sinh(a*t)
--R
--R
--R   (12)  - cos(a t)sinh(a t) + cosh(a t)sin(a t)
--R                                                    Type: Expression(Integer)
--E 12
--S 13 of 27

laplace(%, t, s)
--R
--R
--R              3
--R            4a
--R   (13)  --------
--R          4     4
--R         s  + 4a
--R                                                    Type: Expression(Integer)
--E 13
--S 14 of 27
(exp(a*t) - exp(b*t))/t
--R
--R
--R             b t     a t
--R         - %e    + %e
--R   (14)  ---------------
--R                t
--R                                                    Type: Expression(Integer)
--E 14
--S 15 of 27
laplace(%, t, s)
--R
--R
--R   (15)  - log(s - a) + log(s - b)
--R                                                    Type: Expression(Integer)
--E 15
--S 16 of 27
2/t * (1 - cosh(a*t))
--R
--R
--R         - 2cosh(a t) + 2
--R   (16)  ----------------
--R                 t
--R                                                    Type: Expression(Integer)
--E 16
--S 17 of 27
laplace(%, t, s)
--R
--R
--R              2    2
--R   (17)  log(s  - a ) - 2log(s)
--R                                                    Type: Expression(Integer)
--E 17
--S 18 of 27

2/t * (1 - cos(a*t))
--R
--R
--R         - 2cos(a t) + 2
--R   (18)  ---------------
--R                t
--R                                                    Type: Expression(Integer)
--E 18
--S 19 of 27
laplace(%, t, s)
--R
--R
--R              2    2
--R   (19)  log(s  + a ) - 2log(s)
--R                                                    Type: Expression(Integer)
--E 19
--S 20 of 27
(cos(a*t) - cos(b*t))/t
--R
--R
--R         - cos(b t) + cos(a t)
--R   (20)  ---------------------
--R                   t
--R                                                    Type: Expression(Integer)
--E 20
--S 21 of 27
laplace(%, t, s)
--R
--R
--R              2    2         2    2
--R         log(s  + b ) - log(s  + a )
--R   (21)  ---------------------------
--R                      2
--R                                                    Type: Expression(Integer)
--E 21

———-

We handle some very limited form of special functions — * —

--S 22 of 27
a*Ci(b*t) + c*Si(d*t)
--R
--R
--R   (22)  c Si(d t) + a Ci(b t)
--R                                                    Type: Expression(Integer)

--E 22
--S 23 of 27
laplace(%, t, s)
--R
--R
--R                2    2
--R               s  + b             d
--R         a log(-------) + 2c atan(-)
--R                   2              s
--R                  b
--R   (23)  ---------------------------
--R                      2s
--R                                                    Type: Expression(Integer)
--E 23
--S 24 of 27
exp(a*t+b)*Ei(c*t)
--R
--R
--R                  a t + b
--R   (24)  Ei(c t)%e
--R                                                    Type: Expression(Integer)
--E 24
--S 25 of 27
laplace(%, t, s)
--R
--R
--R           b    s + c - a
--R         %e log(---------)
--R                    c
--R   (25)  -----------------
--R               s - a
--R                                                    Type: Expression(Integer)
--E 25

———-

We keep unknown transforms as formal transform in the answer — * —

--S 26 of 27
sin(a*t) - a*t*cos(a*t) + exp(t**2)
--R
--R
--R                       2
--R                      t
--R   (26)  sin(a t) + %e   - a t cos(a t)
--R                                                    Type: Expression(Integer)

--E 26
--S 27 of 27
laplace(%, t, s)
--R
--R
--R                                     2
--R           4     2 2    4           t           3
--R         (s  + 2a s  + a )laplace(%e  ,t,s) + 2a
--R   (27)  ----------------------------------------
--R                       4     2 2    4
--R                      s  + 2a s  + a
--R                                                    Type: Expression(Integer)
--E 27
)spool
)lisp (bye)

———-

References^{[1] nothing}

10